OpenStudy (anonymous):

Need Some Help, Number Line Probability: A point is chosen at random on the number line between 0 and 1, and the point is colored green. Then another point is chosen at random on the number line between 0 and 1, and this point is colored purple, what is the probability that the number of the purple point is greater than the number of the green point, but less than twice the number of the green point?

OpenStudy (anonymous):

good one... what is the distribution of the green point?

OpenStudy (anonymous):

I'm not quite sure what you are asking

OpenStudy (anonymous):

you need to think about what the underlying distribution (probability distribution) of the green point is in order to get a handle on the probability you seek.

OpenStudy (anonymous):

have you heard of the uniform distribution?

OpenStudy (anonymous):

also, have you heard of conditional probabilities?

OpenStudy (anonymous):

No, I haven't, I might have learned it, just not under that name.

OpenStudy (anonymous):

is this a calculus based course?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Although my course does provide a solution after you correctly answer the problem.

OpenStudy (anonymous):

\[X \text{ ~ } U \left( a, b \right) => P \left( X<x \right)=\frac{ 1 }{ x-a } \text{ for } a<x<b\] this is the continuous version of the uniform distribution which is what you'd need to do this probability. You don't necessarily need calculus to do it but you do need to know about conditional probability. It's difficult to write in this equation editor... let me see if I can work something up for you.

OpenStudy (anonymous):

Hmm, I'm beginning to understand that part. For a, we could put the green point, x, the purple point, and b 2 times the green point... unless I am missing something.

OpenStudy (anonymous):

you're getting there... are the events independent?

OpenStudy (anonymous):

Hmm, I believe so, I'm not sure though

OpenStudy (anonymous):

yes, they are. why, beacuse no restrictions are imposed once we pick our green point. I mean, the purple point can still be anywhere from 0 to 1. that means P( G and P) = P(G) * P(P) but we should still condition on what we get for green. if green is less than .5 then we're looking for P( g<P<2g) and if green is >= .5 then we want P(P>g) = 1-P(G<g)

OpenStudy (anonymous):

What is the difference between G and g? (same with P and p, but not P as in probability? Otherwise, I understand it

OpenStudy (anonymous):

@pgpilot326 If the events are independent then the probability is not conditional....

OpenStudy (anonymous):

Unless you're referring to the fact that if the green point is chosen on a higher number, then the probability of the purple point being greater automatically reduces thus conditional...

OpenStudy (anonymous):

the caps are the random variable the smalls are the values they take on yes, right @genius12 but if we condition, it's much easier to get a handle on what's needed

OpenStudy (anonymous):

in that first one P(g<P<2g), g represents that we got an actual number for G. For example, let's say g = .33, then we can compute P(g<P<2g) = P(.33<P<.66) = (.66 - .33) The distribution I gave earlier is erroneous and I apologize. It should be the length of the interval you want over the total length of the interval. In this problem the total interval is 1. I apologize.

OpenStudy (anonymous):

Ah, okay.

OpenStudy (anonymous):

Hmm, I think i've found a way, using graphing

OpenStudy (anonymous):

Because if x = green point, y = purple point

OpenStudy (anonymous):

0 <= x <= 1, 0 <= y <= 1

OpenStudy (anonymous):



OpenStudy (anonymous):

Then, you would graph the inequality x <= y <= 2x

OpenStudy (anonymous):